OpenStudy (anonymous):

Solve for y. 3 (y+p) = p a. p= -2/3y b p= -3/2y c. y= -2/3p d. y= -3/2p

OpenStudy (anonymous):

\[a(b + c) = ab + ac\] So \(3(y+p) = p\) is the same as \(3y + 3p = p\)

OpenStudy (anonymous):

so what one

OpenStudy (anonymous):

I told you how to figure it out. I'm not going to just tell you the answer.

OpenStudy (anonymous):

how does that formula relate to the awnsers their in fraction form

OpenStudy (anonymous):

Because it's not done. You have to solve for y. If you had \(3y + 3p = p\) how would you get \(y\) by itself?

OpenStudy (anonymous):

by subtracting 3?

OpenStudy (anonymous):

That would give \(3y + 3p - 3 = p - 3\) which doesn't help us much.

OpenStudy (anonymous):

soo what do you do

OpenStudy (anonymous):

Try again!

OpenStudy (anonymous):

hahaha um subtract p

OpenStudy (anonymous):

That's getting there. \(3y + 2p = 0\)

OpenStudy (anonymous):

Its uh c right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

*bows*

OpenStudy (anonymous):

thanks